# Unified Multi-Modal Interleaved Document Representation for Information Retrieval

### v1

@@ 4.4 Further Analysis and Ablation
5 Conclusion
In this paper, we introduced IDentIfy, a novel IR framework designed to address the limitations of conventional methods that rely on solely textual content of documents and their segmented passages. Specifically, our approach sits on top of recent VLMs, which enables integration and representation of diverse multimodal content (including text, images, and tables) into a unified document representation. Also, unlike previous strategies that segment documents at the passage level, our method merges these segments to maintain the document’s structural coherence, while further introducing a reranking strategy for precise identification of relevant sections. Extensive experiments across various IR datasets demonstrated that IDentIfy consistently outperforms existing baselines, confirming that the interleaved multimodal representation significantly enhances the quality of the document retrieval. We believe IDentIfy represents a crucial step toward more comprehensive and contextually aware IR systems, capable of handling the increasing multimodality of modern information sources.
Reproducibility Statement
Our codes are based on publicly available LLaVA-NeXT (Li et al., 2024). The experimental setup and details can be found in Section 4 and Appendix A. The experiments are conducted with publicly available datasets (Mensink et al., 2023; Chen et al., 2023; Lerner et al., 2022; Kweon et al., 2023). We have included our codes in the supplementary material and will publicly release our code.
Appendix
Organization
The supplementary file is organized as follows: In Appendix A, we explain the implementation details for our experiments. In Appendix B, we outline the limitations of our study.
Appendix A Implementation Details
Table 8: Information retrieval datasets summary.
Dataset Query Modality Target Domain Entities Section ID Train Eval Test Corpus size
Encyclopedic-VQA Text, Text-Image Text Species, Landmarks 17k o 177k 2.2k 3.8k 100k
InfoSeek Text-Image Text Diverse 11k x 209k 23k 74k 500k
ViQuAE Text, Text-Image Text Human 1k o 1.2k 1.2k 1.2k 100k
Open-WikiTable Text Table Table - o 3.3k 0.4k 0.4k 1.8k
Dataset configuration
Table 8 summarizes the key properties of the datasets used in our experiment, including query modality, target item, entity domain, number of entities, and whether a section ID is provided to indicate the section containing the answer. Additionally, we provide the number of samples in the training, evaluation, and test splits, as well as the size of the corpus.
Dataset pre-processing
In our study, we leverage interleaved multimodal content from Wikipedia documents. However, the existing corpora associated with our IR datasets often lack this content, typically only including the first few words of each document. Therefore, we augment the corpora by downloading the HTML file of each Wikipedia document.
Since downloading the complete contents for all documents across datasets is time- and memory-intensive, we preprocess the subsets of each corpus, including documents relevant to queries in the training, evaluation, and test splits, as well as unrelated entity documents.
Appendix B Limitations

### v2

@@ Appendix B Additional Experimental Results
Appendix C Case Studies of Document Retrieval
We conduct case studies to demonstrate the advantages of our approach in document retrieval with textual and multimodal queries. In Figure 5 and Figure 6, we illustrate the instances where our approach, which leverages interleaved multimodal contents (e.g., images, tables, and text) within documents, retrieved correct documents for given queries, while the conventional one, which represents documents using only textual data, retrieved documents that appeared to be relevant but were not actually related to the queries.
In Figure 5, a textual query asks for the name of the park located on the north shore of Foster Reservoir. The conventional approach retrieved a document containing unrelated information about a different reservoir. While this document includes terms such as "Peak District National Park" and "North America farm," which make the document superficially relevant, it fails to answer the query. In contrast, our approach identified the document containing the correct answer to the given query.
The advantages of integrating multimodal content into document representation become more apparent in document retrieval with multimodal queries, as shown in Figure 6. For a query consisting of an image of a town hall in Hanover and a textual question about its designer, both our approach and the conventional one retrieved documents about town halls in Germany. However, our approach pinpointed the exact document about the town hall in Hanover, indicating that Hermann Eggert designed the building. The conventional method retrieved a document about a town hall in Munich, which is somewhat related but not an exact match to the query image or question.
These cases underscore the benefits of leveraging multimodal content in information retrieval. Integrating interleaved multimodal elements, our approach aligns more effectively with the input query, resulting in more accurate and fine-grained retrieval. This superiority is supported by Xu et al. (2024), which highlights that models perform better when prompted with rich multimodal information, enabling them to capture alignments across modalities and enhance the representation of given inputs.
Q: What is the name of the park on the north shore of foster reservoir?
(a) Interleaved Multimodal Document Retrieval ,,,,,,,,,,,,, (b) Text-only Document Retrieval
Figure 5: Retrieved documents across different document formats for document retrieval with a given textual query. (a): A document retrieved when represented leveraging interleaved multimodal contents within documents (ours). (b): A document retrieved when using only textual format
[Uncaptioned image] Q: Who designed this building?
(a) Interleaved Multimodal Document Retrieval ,,,,,,,,,,,,, (b) Text-only Document Retrieval
Figure 6: Retrieved documents across different document formats for document retrieval with a given multimodal query. (a): A document retrieved when represented leveraging interleaved multimodal contents within documents (ours). (b): A document retrieved when using only textual format